# Single-Source Regular Path Querying in Terms of Linear Algebra

### v1

@@ 6. Conclusion and Future Work
In this work, we proposed the single-source RPQ evaluation algorithm, which is based on simultaneous traversal of the input graph and the finite automaton specifying path constraints. The traversal is expressed in terms of operations over matrices and vectors, that allows us to provide highly-parallel implementation based on SuiteSparse:GraphBLAS. Our evaluation shows that the suggested algorithm is suitable for in-memory processing of real-world large knowledge graphs. While for some queries the proposed algorithm is slower than competitors, we can conclude that our solution is faster for hard queries: it fits with 1 minute time while other solutions do not.
First of all, it is necessary to analyse abilities to apply well-known optimizations from both RPQ evaluation and BFS algorithms. For example, rare labels (Koschmieder and Leser, 2012) utilization, or push-pull optimization (Yang et al., 2022) respectively.
While it was shown that multiple source BFS can be expressed in terms of linear algebra (Elekes et al., 2020), there is a room for technical optimizations and careful evaluation of the respective modifications to the proposed algorithm, and it should be done in the future.
Thanks to linear algebra, having a single schema of algorithm one can solve different problems varying semiring-like structures. For example, one can look at variations of BFS (Brock et al., 2021) where one can compute reachability facts or information about paths depending on used semiring. In the case of RPQ, there are a number of possible outputs and semantics (Angles et al., 2017): reachability, single path, all paths, simple paths, etc. It is an open question, which of them can be expressed without algorithm changes, but by providing other semiring, and which can be expressed with algorithm changes.
Utilization of GPGPUs to evaluate linear algebra based algorithms for graph analysis can significantly improve performance in some cases (Yang et al., 2022; Stanislavovic, 2023). It is necessary to investigate, whether utilization of GPGPU in our algorithm improve performance or not.
Distributed solutions are a way for huge graphs processing (Guo et al., 2021b). Implementation and evaluation of the proposed algorithm in distributed settings, for example, using CombBLAS (Buluç and Gilbert, 2011) that provides distributed linear algebraic routines for graph analysis, is also a task for the future.
@@ Acknowledgements.
This research was supported by the St Petersburg University, grant id 116636233.
@@ Appendix A Proof of Correctness
The proposed algorithm, represented in 1, computes matrix $P$ such that respective relation $\mathcal{P}\subseteq Q\times V$ has the following property.
@@
Label Count
P31 95 749 143
P17 13 994 387
P131 10 609 188
P106 8 533 132
P279 3 128 156
P19 2 800 454
P40 1 419 672
P39 1 302 690
P495 1 272 895
P20 1 127 738
P138 322 330
P551 237 546
Table 2. Number of edges with specific labels used in queries
Query Speedup #Answer
1 Q13442814 ^P31 ?x1 19.09 37 280 216
2 ?x1 (P17 / P31) Q6256 14.49 12 360 733
3 ?x1 (P31* / P279*)+ Q268592 13.34 60 480 400
4 ?x1 ((P106 | P39 | P31) / P279*) Q488383 12.79 76 388 701
5 ?x1 (P31 / P279*) Q191067 11.72 38 410 384
6 ?x1 (P31 / P279*) Q488383 11.10 75 238 748
7 ?x1 (P31 / P279*) Q35120 10.50 90 321 703
8 ?x1 (P31 / P279*) Q13442814 10.44 37 325 506
9 ?x1 (P31 / P279*) Q386724 9.71 55 722 209
10 ?x1 (P31 / P279*) Q7184903 9.53 58 080 849
11 Q10281806 P131* ?x1 0.16 6
12 Q1628288 P131* ?x1 0.17 6
13 Q111 (P31 / P279*) ?x1 0.19 18
14 Q939035 P131* ?x1 0.19 6
15 Q1549583 P131* ?x1 0.20 6
16 Q481075 P131* ?x1 0.20 5
17 Q29873614 P131* ?x1 0.21 4
18 ?x1 P40+ Q748955 0.21 17 699
19 Q28733811 P131* ?x1 0.21 5
20 ?x1 P40+ Q390192 0.23 22 335
Table 3. Top 10 relatively slowest and relatively fastest queries for MDB
Denote state of matrix $M$ after the step $n$ of the loop on lines 7–10 of the algorithm 1 as $M_{n}$ and state of $P$ as $P_{n}$. Introduce auxiliary relations $\mathcal{M}_{n}\subseteq Q\times V$ represented by $M_{n}$ and $\mathcal{P}_{n}\subseteq Q\times V$ represented by $P$ after the step $n$. $\mathcal{P}$ and $\mathcal{P}_{n}$ are connected with relations $\mathcal{M}_{n}$:
@@
The approach is to build a relation between the states of automaton and the vertices of the graph and update it by traversing both automaton and the graph at the same time. As an algorithm invariant after the step $n$ we claim this properties for $\mathcal{M}_{n}$ represented by the matrix $M_{n}$:
@@
Note, that if we continue to loop forever in lines 7–10 from $\mathcal{M}_{n}=\operatorname{\varnothing}$ it follows $\mathcal{M}_{m}=\operatorname{\varnothing}$ for $m>n$. It means, it’s enough to iterate until the $M$ matrix becomes $\operatorname{\operatorname{\mathbf{0}}}$.
Notice, that $\mathcal{P}_{n}\subsetneq\mathcal{P}_{n+1}$ or $|\mathcal{P}_{n}|<|\mathcal{P}_{n+1}|$ if $\mathcal{M}_{n+1}$ is not empty. And $\mathcal{M}_{n+1}\cap\mathcal{P}_{n}=\operatorname{\varnothing}$ for all $n\in\mathbb{N}$. $|\mathcal{P}_{n}|\leq|Q||V|$ and $|\mathcal{M}_{n}|\leq|Q||V|$ for all $n\in\mathbb{N}$. Thus, if $\mathcal{M}_{m}\neq\operatorname{\varnothing}$ for all $m<|Q||V|$ then $\mathcal{M}_{|Q||V|}=\operatorname{\varnothing}$, since $|\mathcal{P}_{|Q||V|-1}|\geq|Q||V|$. This means the algorithm always finishes in maximum of $|Q||V|$ steps.
Obviously, the invariant is held for $n=0$ after the initialization of the matrix $M$ with $M_{0}$. If you consider the paths of length 00 the set of vertices and automaton states coincides with $\{v_{s}\}$ and $Q_{S}$ correspondingly.
@@
Consider evaluating the $n+1$ step of the algorithm. Fix the label $a\in L$. After the step $n$ $M$ represents a relation $\mathcal{M}_{n}$. At first, evaluate the first matrix product $M^{\prime a}_{n+1}=(N^{a})^{T}\operatorname{\operatorname{\otimes}}M$. This product represents a relation $\mathcal{M}^{\prime a}_{n+1}\subseteq Q\times V$:
@@
The updating the relation between the 2-NFA states and the graph vertices $\mathcal{M}$ with $\mathcal{M}^{\prime}$ value is follows form the following properties.
@@
Ensure the invariant is preserved for $(q^{\prime},v^{\prime})$ in $\mathcal{M}^{\prime}$. Let $\pi_{\mathcal{G}}=(e_{1},...,e_{n})$, $\pi_{\mathcal{N}}=(\delta_{1},...,\delta_{n})$ the paths to $(q,v)$ satisfying the conditions A. Then:
@@
  • $\mathcal{P}_{n}=\displaystyle\bigcup_{m\leq n}\mathcal{M}_{m}$, thus, if $(q,v)\in\mathcal{M}_{m}$ for $m\leq n$ then $(q,v)\notin\mathcal{M}_{n+1}$.
Since, the conditions (A) are preserved for the relation $\mathcal{M}_{n+1}$ matching the new value of $M\leftarrow M_{n+1}$ the invariant is preserved.
Appendix B Queries Statistics
For detailed analysis of performance we collected top 10 relatively fastest and relatively slowest queries for each system. Results presented in tables 345. Queries represented in SPARQL-like syntax. Speedup means speeding up of the proposed solution relative to respective system. #Answer is a number of reachable vertices. Also we collected total number of edges in the graph with specific labels appeared in collected queries. Results presented in tables 2.
Query Speedup #Answer
1 Q7187 ((P279 / (P279* | P31*)) | (P31/(P279* | P31*))) ?x1 506 771.72 18
2 Q5 ((P279 / (P279* | P31*)) | (P31 / (P279* | P31*))) ?x1 345 401.17 30
3 Q11629 ((P279 / (P279* | P31*)) | (P31 / (P279* | P31*))) ?x1 311 512.45 63
4 Q1656682 ((P279 | P31) / (P279* | P31*)) ?x1 306 341.47 7
5 Q5 (P279 / (P279* | P31*)) ?x1 241 437.89 20
6 Q5 ((P279 | P31) / (P279* | P31*)) ?x1 185 795.87 30
7 Q81328 (P31 / P279? / P279? / P279? / P279? / P279?) ?x1 127 813.38 18
8 Q4235 (P31 / P279? / P279? / P279? / P279? / P279?) ?x1 123 845.99 18
9 Q2 (P31/P279? / P279? / P279? / P279? / P279? / P279?) ?x1 122 229.53 20
10 Q111 (P31 / P279? / P279? / P279? / P279? / P279? / P279?) ?x1 92 952.34 11
11 ?x1 (P17 | P495) Q30 0.03 1 697 681
12 ?x1 (P138 / P106) Q1028181 0.04 8 135
13 ?x1 (P31 / P279) Q18918145 0.05 188 351
14 ?x1 (P19 | P20 | P551) Q90 0.06 50 972
15 ?x1 (P31 / P279) Q16521 0.07 245 925
16 ?x1 (P19 | P20 | P551) Q1741 0.07 43 405
17 ?x1 (P31 / P279*) Q18918145 0.09 217 394
18 ?x1 P31+ Q4830453 0.09 190 623
19 ?x1 (P31 / P279) Q1030034 0.09 82 052
20 ?x1 (P31 / P279) Q486972 0.09 581 652
Table 4. Top 10 relatively slowest and relatively fastest queries for RPQ-Matrix
Query Speedup #Answer
1 ?x1 (P31 / P279*) Q959309 37327.83 1605
2 ?x1 (P31 / P279*) Q385337 30592.48 1980
3 ?x1 (P31/P279*) Q543654 10783.13 7318
4 Q64 (P31/P279*) Q811430 8052.45 115
5 ?x1 P279* Q16707842 2000.93 59
6 ?x1 P279* Q15630906 1664.48 13
7 ?x1 P131* Q168159 1335.28 267
8 ?x1 P279* Q144 606.69 269
9 ?x1 P131* Q108053 598.19 2137
10 ?x1 (P31 / P279*) Q1867183 596.26 1217
11 ?x1 (P138 / P106) Q1028181 0.66 8 135
12 Q257023 P131* ?x1 1.00 6
13 Q326313 P131* ?x1 1.19 6
14 Q204756 P131* ?x1 1.23 5
15 Q29018634 P131* ?x1 1.26 7
16 Q24192592 P131* ?x1 1.26 7
17 Q5964731 P131* ?x1 1.29 6
18 Q328046 P131* ?x1 1.30 6
19 Q2321874 P131* ?x1 1.33 5
20 Q894219 P131* ?x1 1.34 5
Table 5. Top 10 relatively slowest and relatively fastest queries for FalkorDB

### v2

@@ 6. Conclusion and Future Work
In this work, we proposed the single-source RPQ evaluation algorithm, which is based on the simultaneous traversal of the input graph and the finite automaton specifying path constraints. The traversal is expressed in terms of operations over matrices and vectors, which allows us to provide a highly parallel implementation based on SuiteSparse:GraphBLAS.
Our experimental evaluation shows that the suggested algorithm is suitable for in-memory processing of real-world large knowledge graphs. While for some queries the proposed algorithm is slower than competitors, we can conclude that our solution is faster for hard queries: it fits with 1 minute time while other solutions do not.
Our results demonstrate that both algorithmic design and implementation details of underlying linear algebra primitives significantly impact 2-RPQ performance. While average performance metrics provide straightforward comparisons, specific cases require deeper analysis to identify each algorithm’s strengths under different conditions; to understand implementation trade-offs; to guide development of robust universal solutions.
Furthermore, this analysis provides valuable insights for enhancing the GraphBLAS API by identifying critical functionality required for efficient RPQ evaluation algorithms. The performance characteristics we observed highlight specific linear algebra primitives that most significantly impact query processing efficiency, suggesting potential directions for API optimization.
Regarding our new algorithm, first of all, it is necessary to analyse abilities to apply well-known optimizations from both RPQ evaluation and BFS algorithms. For example, rare labels (Koschmieder and Leser, 2012) utilization, or push-pull optimization (Yang et al., 2022) respectively.
Although multiple source BFS has been shown to be expressed in terms of linear algebra (Elekes et al., 2020), there is room for technical optimizations and careful evaluation of the respective modifications to the proposed algorithm, and it should be done in the future.
Thanks to linear algebra, having a single schema of algorithm one can solve different problems varying semiring-like structures. For example, one can look at variations of the BFS (Brock et al., 2021) where one can compute reachability facts or information about paths depending on the used semiring. In the case of RPQ, there are a number of possible outputs and semantics (Angles et al., 2017): reachability, single path, all paths, simple paths, etc. It is an open question, which of them can be expressed without algorithm changes, but by providing other semirings, and which can be expressed with algorithm changes.
Utilization of GPGPUs to evaluate linear algebra-based algorithms for graph analysis can significantly improve performance in some cases (Yang et al., 2022; Stanislavovic, 2023). It is necessary to investigate, whether utilization of GPGPU in our algorithm improves performance or not.
Distributed solutions are a way to process graph processing (Guo et al., 2021b). Implementation and evaluation of the proposed algorithm in distributed settings, for example, using CombBLAS (Buluç and Gilbert, 2011) that provides distributed linear algebraic routines for graph analysis, is also a task for the future.
@@ Acknowledgements.
This research has been supported by the St. Petersburg State University, grant id 116636233. Paper submission has been sponsored by Tarantool DBMS team, VK Tech LLC.
@@ Appendix A Proof of Correctness
The proposed algorithm, represented in 1, computes the matrix $P$ such that the respective relation $\mathcal{P}\subseteq Q\times V$ has the following property.
@@
Denote the state of matrix $M$ after the step $n$ of the loop on lines 7–10 of algorithm 1 as $M_{n}$ and the state of $P$ as $P_{n}$. Introduce the auxiliary relations $\mathcal{M}_{n}\subseteq Q\times V$ represented by $M_{n}$ and $\mathcal{P}_{n}\subseteq Q\times V$ represented by $P$ after step $n$. $\mathcal{P}$ and $\mathcal{P}_{n}$ are connected with relations $\mathcal{M}_{n}$:
@@
The approach is to build a relation between the automaton states and the vertices of the graph and update it by traversing both automaton and graph at the same time. As an algorithm invariant after the step $n$ we claim this properties for $\mathcal{M}_{n}$ represented by the matrix $M_{n}$:
@@
Note that if we continue looping forever in lines 7–10 from $\mathcal{M}_{n}=\operatorname{\varnothing}$ it follows $\mathcal{M}_{m}=\operatorname{\varnothing}$ for $m>n$. This means that it is enough to iterate until the $M$ matrix becomes $\operatorname{\operatorname{\mathbf{0}}}$.
Notice that $\mathcal{P}_{n}\subsetneq\mathcal{P}_{n+1}$ or $|\mathcal{P}_{n}|<|\mathcal{P}_{n+1}|$ if $\mathcal{M}_{n+1}$ is not empty. And $\mathcal{M}_{n+1}\cap\mathcal{P}_{n}=\operatorname{\varnothing}$ for all $n\in\mathbb{N}$. $|\mathcal{P}_{n}|\leq|Q||V|$ and $|\mathcal{M}_{n}|\leq|Q||V|$ for all $n\in\mathbb{N}$. Thus, if $\mathcal{M}_{m}\neq\operatorname{\varnothing}$ for all $m<|Q||V|$ then $\mathcal{M}_{|Q||V|}=\operatorname{\varnothing}$, since $|\mathcal{P}_{|Q||V|-1}|\geq|Q||V|$. This means that the algorithm always finishes in a maximum of $|Q||V|$ steps.
Obviously, the invariant holds for $n=0$ after initializing the matrix $M$ with $M_{0}$. If you consider the paths of length 00, the set of vertices and automaton states coincides with $\{v_{s}\}$ and $Q_{S}$ correspondingly.
@@
Consider evaluating the $n+1$ step of the algorithm. Fix the label $a\in L$. After step $n$, $M$ represents a relation $\mathcal{M}_{n}$. At first, evaluate the first matrix product $M^{\prime a}_{n+1}=(N^{a})^{T}\operatorname{\operatorname{\otimes}}M$. This product represents a relation $\mathcal{M}^{\prime a}_{n+1}\subseteq Q\times V$:
@@
Updating the relation between the 2-NFA states and the graph vertices $\mathcal{M}$ with the $\mathcal{M}^{\prime}$ value derives the following properties.
@@
Ensure that the invariant is preserved for $(q^{\prime},v^{\prime})$ in $\mathcal{M}^{\prime}$. Let $\pi_{\mathcal{G}}=(e_{1},...,e_{n})$, $\pi_{\mathcal{N}}=(\delta_{1},...,\delta_{n})$ be the paths to $(q,v)$ satisfying conditions A. Then:
@@
  • $\mathcal{P}_{n}=\displaystyle\bigcup_{m\leq n}\mathcal{M}_{m}$, so if $(q,v)\in\mathcal{M}_{m}$ for $m\leq n$ then $(q,v)\notin\mathcal{M}_{n+1}$.
Since conditions (A) are preserved for the relation $\mathcal{M}_{n+1}$ matching the new value of $M\leftarrow M_{n+1}$, the invariant is preserved.